THE PERCENTAGE OF IONIC CHARACTER OF A COMPOUND HAVING 1.2d DIPOLE MOMENT AND 125 PM BOND LENGTH IS [ANS:20 PERCENT}

Grade:11

1 Answers

Ssv goyal
5 years ago
hello there, i got the sol.
ionic character of a compound =observed dipole moment/calculated dipole moment
value of d=3.3*10-30  and 1pm=10-12 m.
calculated dipole moment=Q*separation
e*125*10-12
multiply by100 to get the %